OpenStudy (anonymous):

Please help. 80 = 113+65 - (2√ (113 x 65)) x cosA cosA = 0.5717 A = 55 How did they get to step 2?

OpenStudy (anonymous):

80 - 113 - 65 = -2 root (113*65) cos A -98 = -2 root(7345) cos A -98/ (-2root(7345) = cosA Evaluate... -0.5717

OpenStudy (anonymous):

Thanks!